# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

### v1

@@ 3.1.3   Extension of the approximate solution onto the whole time interval $[0,T]$
In Step 1, we constructed the coefficients $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ as a solution of a nonlinear system of ODEs existing on its right-maximal time interval $[0,T_{k}^{*})\cap[0,T]$. We now assume that $T_{k}^{*}\leq T$. By the definition of the approximate solutions given in (3.3) and the uniform bound (3.1.2), we infer that for any $T_{k}\in[0,T_{k}^{*})$, all $t\in[0,T_{k}]$, and all $i\in\{1,...,k\}$,
@@
This means that the solution $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ is bounded on the time interval $[0,T_{k}^{*})$ by a constant that is independent of $T_{k}$ and $k$. Hence, according to classical ODE theory, the solution can thus be extended beyond the time $T_{k}^{*}$. However, as the solution was assumed to be right-maximal, this is a contradiction. We thus have $T_{k}^{*}>T$, which directly implies $[0,T_{k}^{*})\cap[0,T]=[0,T]$. This means that the solution $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ of the ODE system actually exists on the whole time interval $[0,T]$. As the coefficients $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ can be reconstructed from $\mathbf{a}^{k}$ and $\mathbf{b}^{k}$ by the corresponding system of algebraic equations, they also exist on the whole time interval $[0,T]$. Recalling (3.3) and the construction of ${\boldsymbol{v}}_{k}$ this directly entails that the approximate solution $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ actually exists in $[0,T]$. Hence, choosing $T_{k}=T$ in (3.1.2), we eventually conclude
@@ 3.1.4   Convergence to a weak solution as $k\to\infty$
${\boldsymbol{v}}_{k}$ $\to{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$},$ (3.21a)
${\boldsymbol{v}}_{k}|_{\Gamma}$ $\to{\boldsymbol{v}}|_{\Gamma}$ weakly in $\quad\text{weakly in $L^{2}(0,T;\mathbf{H}_{\Gamma})$},$ (3.21b)
[…]
strongly in C0([0,T];Hs(Ω)), and a.e. in $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~{}$Q$}},$ (3.21c)
[…]
strongly in C0([0,T];Hs(Γ)), and a.e. on $\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~{}$\Sigma$}},$ (3.21d)
$\mu_{k}$ $\to\mu$ weakly in $\quad\text{weakly in $L^{2}(0,T;V)\cap L^{4}(0,T;H)$},$ (3.21e)
$\theta_{k}$ $\to\theta$ weakly in $\quad\text{weakly in $L^{2}(0,T;{V}_{\Gamma})\cap L^{4}(0,T;{H}_{\Gamma})$},$ (3.21f)
@@
Due to the trace theorem, the strong convergence from (3.1.4) (with $s>\frac{1}{2}$) directly yields
@@
$\left\|F^{\prime}(\varphi_{k})\right\|_{L^{\frac{6}{5}}(Q)}\leq C\quad\text{and}\quad\left\|G^{\prime}(\varphi_{k})\right\|_{L^{2}(\Sigma)}\leq C.$
Hence, there exist weakly convergent subsequences in the respective spaces. As $F^{\prime}$ and $G^{\prime}$ are continuous, we use the pointwise convergences from (3.1.4) and (3.1.4) to conclude
$F^{\prime}(\varphi_{k})$ $\to F^{\prime}(\varphi)$ weakly in L65(Q) and a.e. in $\quad\text{weakly in $L^{\frac{6}{5}}(Q)$ and {a.e.\ in~{}$Q$}},$ (3.23)
@@
since the weak limit and the pointwise limit must coincide (see, e.g., [26, Proposition 9.2c]). Furthermore, it follows from the pointwise convergences in (3.1.4) and (3.1.4) that, as $k\to\infty$,
[…]
@@
as the functions $M_{\Omega}$, $M_{\Gamma}$, $\nu$, $\lambda$ and $\gamma$ are continuous. Since, due to (A2) and (A3), these functions are also bounded, we use Lebesgue’s dominated convergence theorem along with the weak convergences in (3.1.4) to infer that
$\nu(\varphi_{k})\mathbf{D}{\boldsymbol{v}}_{k}$ $\to\nu(\varphi)\mathbf{D}{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d\times d})$},$ (3.27)
$\lambda(\varphi_{k}){\boldsymbol{v}}_{k}$ $\to\lambda(\varphi){\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.28)
$M_{\Omega}(\varphi_{k})\nabla\mu_{k}$ $\to M_{\Omega}(\varphi)\nabla\mu$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.29)
$\gamma(\psi_{k}){\boldsymbol{v}}_{k}$ $\to\gamma(\psi){\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$},$ (3.30)
$M_{\Gamma}(\psi_{k})\nabla_{\Gamma}\theta_{k}$ $\to M_{\Gamma}(\psi)\nabla_{\Gamma}\theta$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$}.$ (3.31)
Combining the convergences (3.21a)–(3.21f), (3.22)–(3.24) and (3.27)–(3.31), it is straightforward to pass to the limit as $k\to\infty$ in the discretized weak formulation (3.1.1) to conclude that the quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ fulfills the variational formulation ((ii)) for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$,
@@
Hence, because of density, ((ii)) holds true for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V$, ${\zeta_{\Gamma}}\in V_{\Gamma}$ and $(\eta,{\eta_{\Gamma}})\in{\cal V}={\cal V}_{K}$. This verifies Definition 2.3(ii).
@@
as the orthogonal projections converge strongly in $H$ and in ${H}_{\Gamma}$, respectively. On the other hand, it follows from the strong convergences in (3.1.4) and (3.1.4) that
@@
We still need to establish the weak energy dissipation law. Therefore, let $\rho\in C^{\infty}([0,T])$ be an arbitrary nonnegative test function. Employing the convergences (3.1.4) and (3.1.4), the weak lower semicontinuity of the mappings
@@
$\sqrt{\nu(\varphi_{k})}\,\mathbf{D}{\boldsymbol{v}}_{k}$ $\to\sqrt{\nu(\varphi)}\,\mathbf{D}{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d\times d})$},$ (3.33)
$\sqrt{\lambda(\varphi_{k})}\,{\boldsymbol{v}}_{k}$ $\to\sqrt{\lambda(\varphi)}\,{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.34)
$\sqrt{M_{\Omega}(\varphi_{k})}\,\nabla\mu_{k}$ $\to\sqrt{M_{\Omega}(\varphi)}\,\nabla\mu$ weakly in $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.35)
$\sqrt{\gamma(\psi_{k})}\,{\boldsymbol{v}}_{k}$ $\to\sqrt{\gamma(\psi)}\,{\boldsymbol{v}}$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$},$ (3.36)
$\sqrt{M_{\Gamma}(\psi_{k})}\,\nabla_{\Gamma}\theta_{k}$ $\to\sqrt{M_{\Gamma}(\psi)}\,\nabla_{\Gamma}\theta$ weakly in $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$}.$ (3.37)
@@
Here, invoking the growth conditions from (R1), the equality in the last line follows by means of Lebesgue’s general convergence theorem (see [8, Section 3.25]) since the orthogonal projections in (3.5) converge strongly in $V$ and in ${V}_{\Gamma}$, respectively. As the nonnegative test function $\rho$ was arbitrary, this proves that the weak energy dissipation law stated in ((iv)) holds for almost all $t\in[0,T]$. As the time integral in this inequality is continuous with respect to $t$ and since the mapping $t\mapsto E_{K}\big{(}\varphi(t),\psi(t)\big{)}$ is lower semicontinuous, we conclude that ((iv)) actually holds true for all $t\in[0,T]$. This means that Definition 2.3(iv) is verified.
@@ 3.1.5   Additional regularity for the phase-fields
To prove additional regularity for the phase-fields, we assume $\Omega$ to be of class $C^{\ell}$ with $\ell\in\{2,3\}$, and we return to the Faedo–Galerkin scheme. First of all, applying regularity theory for Poisson’s equation with an inhomogeneous Neumann boundary condition (see, e.g., [52, Theorem 4.18] or [59, s. 5, Proposition 7.7]), we infer that the eigenfunctions $u_{i}$ exhibit the regularity $u_{i}\in H^{\ell}(\Omega)$ for all $i\in{\mathbb{N}}$. Moreover, applying regularity theory for elliptic equations on submanifolds (see, e.g., [59, s. 5, Theorem 1.3]), we conclude that the eigenfunctions $v_{i}$ have the regularity $v_{i}\in H^{\ell}(\Gamma)$ for all $i\in{\mathbb{N}}$. This directly entails
$\varphi_{k}(t)\in V_{k}\cap H^{\ell}(\Omega)\quad\text{and}\quad\psi_{k}(t)\in V_{\Gamma,k}\cap H^{\ell}(\Gamma)$
for all $t\in[0,T]$.
Let now $t\in[0,T]$ be arbitrary. We infer from (3.1.1) that there exists a null set $\mathcal{N}\subset[0,T]$ such that
$\int_{\Omega}\nabla\varphi_{k}(t)\cdot\nabla\eta+\int_{\Gamma}\nabla_{\Gamma}\psi_{k}(t)\cdot\nabla_{\Gamma}{\eta_{\Gamma}}+\frac{1}{K}\int_{\Gamma}\big{(}\psi_{k}(t)-\varphi_{k}(t)\big{)}({\eta_{\Gamma}}-\eta)$
$\quad=\int_{\Omega}\Big{(}\mu_{k}(t)-\mathbb{P}_{V_{k}}\big{[}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{]}\Big{)}\,\eta+\int_{\Gamma}\Big{(}\theta_{k}(t)-\mathbb{P}_{V_{\Gamma,k}}\big{[}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{]}\Big{)}\,{\eta_{\Gamma}}$ (3.38)
for all test functions $(\eta,{\eta_{\Gamma}})\in{\cal V}$ and all $t\in[0,T]\setminus\mathcal{N}$.
Let now $t\in[0,T]\setminus\mathcal{N}$ be arbitrary. Then, (3.1.5) entails that the pair $\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}$ is a weak solution of the bulk-surface elliptic problem
$-\Delta\varphi_{k}(t)$ $=f_{k}(t)$ in $\quad\text{in $\Omega$},$ (3.39a)
$-\Delta_{\Gamma}\psi_{k}(t)+\partial_{\bf n}\varphi_{k}(t)$ $=g_{k}(t)$ on $\quad\text{on $\Gamma$},$ (3.39b)
$K\partial_{\bf n}\varphi_{k}(t)$ $=\psi_{k}(t)-\varphi_{k}(t)$ on $\quad\text{on $\Gamma$},$ (3.39c)
@@
$f_{k}(t)$ $:=\mu_{k}(t)-\mathbb{P}_{V_{k}}\big{[}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{]}$ $\in V_{k}\subset V,$
$g_{k}(t)$ $:=\theta_{k}(t)-\mathbb{P}_{V_{\Gamma,k}}\big{[}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{]}$ $\in V_{\Gamma,k}\subset{V}_{\Gamma}.$
In the following, without loss of generality, we assume that the growth conditions from (R1) hold with $p\in[5,6)$. Most part of the computations to follow resemble those performed in [40]. Therefore, we just highlight the key steps for the reader’s convenience. Using the uniform estimate (3.1.3), the Gagliardo–Nirenberg inequality and Young’s inequality, we derive the estimate
$\big{\|}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{\|}_{H}$ $\leq C+C\|\varphi_{k}(t)\|_{L^{2(p-1)}(\Omega)}^{p-1}$
$\leq C+C\|\varphi_{k}(t)\|_{L^{6}(\Omega)}^{\frac{p+2}{2}}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{\frac{p-4}{2}}$
$\leq C\varepsilon^{-1}+\varepsilon\|\varphi_{k}(t)\|_{H^{2}(\Omega)},$ (3.40)
for any $\varepsilon\in(0,1)$. Moreover, due to (R1) and (3.1.3), we have
$\big{\|}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{\|}_{{H}_{\Gamma}}\leq C+C\|\psi_{k}(t)\|_{L^{2(q-1)}(\Gamma)}^{q-1}\leq C.$ (3.41)
We first consider the case $\ell=2$. Applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we deduce $\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}\in H^{2}(\Omega)\times H^{2}(\Gamma)$ with
$\left\|\varphi_{k}(t)\right\|_{H^{2}(\Omega)}^{2}+\left\|\psi_{k}(t)\right\|_{H^{2}(\Gamma)}^{2}\leq C\left\|f_{k}(t)\right\|_{H}^{2}+C\left\|g_{k}(t)\right\|_{{H}_{\Gamma}}^{2}$
$\quad\leq C\left\|\mu_{k}(t)\right\|_{H}^{2}+C\left\|\theta_{k}(t)\right\|_{{H}_{\Gamma}}^{2}+C(1+\varepsilon^{-2})+C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{2}.$
After choosing $\varepsilon\in(0,1)$ sufficiently small, we absorb the term $C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{2}$ by the left-hand side. Squaring and integrating the resulting inequality with respect to $t$ from 00 to $T$, we use (3.1.3) to arrive at the uniform estimate
$\left\|\varphi_{k}\right\|_{L^{4}(0,T;H^{2}(\Omega))}+\left\|\psi_{k}\right\|_{L^{4}(0,T;H^{2}(\Gamma))}\leq C.$ (3.42)
Consequently, we have, as $k\to\infty$,
$\varphi_{k}$ $\to\varphi\quad\text{weakly in $L^{4}(0,T;H^{2}(\Omega))$},$
$\psi_{k}$ $\to\psi\quad\text{weakly in $L^{4}(0,T;H^{2}(\Gamma))$},$
after another subsequence extraction. This proves the regularity assertion (2.15a). As a direct consequence, it follows via integration by parts and the fundamental lemma of calculus that the equations (1.1d), (1.1f) and (1.1g) are now even satisfied in the strong sense (i.e., a.e. in $Q$ and on $\Sigma$, respectively).
Let us now consider the case $\ell=3$. Employing the uniform estimate (3.1.3), the Gagliardo–Nirenberg inequality, Agmon’s inequality and Young’s inequality, we derive the estimate
$\big{\|}F^{\prime\prime}\big{(}\varphi_{k}(t)\big{)}\nabla\varphi_{k}(t)\big{\|}_{{\bf H}}$ $\leq C\|\nabla\varphi_{k}(t)\|_{{\bf H}}+C\||\varphi_{k}(t)|^{p-2}\nabla\varphi_{k}(t)\|_{{\bf H}}$
$\leq C+C\|\varphi_{k}(t)\|_{L^{2(p-2)}(\Omega)}^{p-2}\|\nabla\varphi_{k}(t)\|_{{\bf{L}}^{\infty}(\Omega)}$
$\leq C+C\|\varphi_{k}(t)\|_{L^{6}(\Omega)}^{\frac{p+1}{2}}\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{\frac{p-5}{2}+\frac{1}{2}}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{1}{2}}$
$\leq C+C\|\varphi_{k}(t)\|_{H^{2}(\Omega)}^{\frac{p-4}{2}}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{1}{2}}$
$\leq C+C\|\varphi_{k}(t)\|_{H^{1}(\Omega)}^{\frac{p-4}{4}}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{p-4}{4}+\frac{1}{2}}$
$\leq C+C\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{\frac{p-2}{4}}$
$\leq C\varepsilon^{-1}+\varepsilon\|\varphi_{k}(t)\|_{H^{3}(\Omega)}$ (3.43)
for any $\varepsilon\in(0,1)$. In combination with (3.1.5), this proves that
$\big{\|}F^{\prime}\big{(}\varphi_{k}(t)\big{)}\big{\|}_{V}\leq C\varepsilon^{-1}+\varepsilon\|\varphi_{k}(t)\|_{H^{3}(\Omega)}.$ (3.44)
In a similar fashion, we derive the estimate
$\big{\|}G^{\prime}\big{(}\psi_{k}(t)\big{)}\big{\|}_{{V}_{\Gamma}}\leq C\varepsilon^{-1}+\varepsilon\|\psi_{k}(t)\|_{H^{3}(\Gamma)}$ (3.45)

### v2

@@ 3.1.3 Extension of the approximate solution onto the whole time interval $[0,T]$
In Step 1, we constructed the coefficients $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ as a solution of a nonlinear system of ODEs existing on its right-maximal time interval $0,T_{k}^{*})\cap[0,T]$. We now assume that $T_{k}^{*}\leq T$. By the definition of the approximate solutions given in (3.3) and the uniform bound (3.1.2), we infer that for any $T_{k}\in[0,T_{k}^{*})$, all $t\in[0,T_{k}]$, and all $i\in\{1,...,k\}$,
@@
This means that the solution $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ is bounded on the time interval $0,T_{k}^{*})$ by a constant that is independent of $T_{k}$ and $k$. Hence, according to classical ODE theory, the solution can thus be extended beyond the time $T_{k}^{*}$. However, as the solution was assumed to be right-maximal, this is a contradiction. We thus have $T_{k}^{*}>T$, which directly implies $0,T_{k}^{*})\cap[0,T]=[0,T]$. This means that the solution $(\mathbf{a}^{k},\mathbf{b}^{k})^{\top}$ of the ODE system actually exists on the whole time interval $[0,T]$. As the coefficients $\mathbf{c}^{k}$ and $\mathbf{d}^{k}$ can be reconstructed from $\mathbf{a}^{k}$ and $\mathbf{b}^{k}$ by the corresponding system of algebraic equations, they also exist on the whole time interval $[0,T]$. Recalling (3.3) and the construction of ${\boldsymbol{v}}_{k}$ this directly entails that the approximate solution $({\boldsymbol{v}}_{k},\varphi_{k},\mu_{k},\psi_{k},\theta_{k})$ actually exists in $[0,T]$. Hence, choosing $T_{k}=T$ in (3.1.2), we eventually conclude
@@ 3.1.4 Convergence to a weak solution as $k\to\infty$
${\boldsymbol{v}}_{k}$ $\to{\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(0,T;{\bf V}_{\sigma,{\bf n}})$},$ (3.25a)
${\boldsymbol{v}}_{k}|_{\Gamma}$ $\to{\boldsymbol{v}}|_{\Gamma}$ $\quad\text{weakly in $L^{2}(0,T;\mathbf{H}_{\Gamma})$},$ (3.25b)
$\varphi_{k}$ $\to\varphi$ $\quad\text{weakly-${}^{*}$ in $L^{\infty}(0,T;V)$, weakly in $H^{1}(0,T;V^{*})$},$
$\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Omega))$, and {a.e.\ in~$Q$}},$ (3.25c)
$\psi_{k}$ $\to\psi$ $\quad\text{weakly-${}^{*}$ in $L^{\infty}(0,T;{V}_{\Gamma})$, weakly in $H^{1}(0,T;{V}_{\Gamma}^{*})$},$
$\qquad\text{strongly in $C^{0}([0,T];H^{s}(\Gamma))$, and {a.e.\ on~$\Sigma$}},$ (3.25d)
$\mu_{k}$ $\to\mu$ $\quad\text{weakly in $L^{2}(0,T;V)\cap L^{4}(0,T;H)$},$ (3.25e)
$\theta_{k}$ $\to\theta$ $\quad\text{weakly in $L^{2}(0,T;{V}_{\Gamma})\cap L^{4}(0,T;{H}_{\Gamma})$},$ (3.25f)
@@
Due to the trace theorem, the strong convergence from (3.25) (with $s>\frac{1}{2}$) directly yields
@@
$\left\|F^{\prime}(\varphi_{k})\right\|_{L^{6/5}(Q)}\leq C\quad\text{and}\quad\left\|G^{\prime}(\varphi_{k})\right\|_{L^{2}(\Sigma)}\leq C.$
Hence, there exist weakly convergent subsequences in the respective spaces. As $F^{\prime}$ and $G^{\prime}$ are continuous, we use the pointwise convergences from (3.25) and (3.25) to conclude
$F^{\prime}(\varphi_{k})$ $\to F^{\prime}(\varphi)$ $\quad\text{weakly in $L^{6/5}(Q)$ and {a.e.\ in~$Q$}},$ (3.27)
@@
since the weak limit and the pointwise limit must coincide (see, e.g., [32, Proposition 9.2c]). Furthermore, it follows from the pointwise convergences in (3.25) and (3.25) that, as $k\to\infty$,
$M_{\Omega}(\varphi_{k})$ $\to M_{\Omega}(\varphi),$ $\quad\nu(\varphi_{k})$ $\to\nu(\varphi),\quad\lambda(\varphi_{k})\to\lambda(\varphi)$ $\quad\text{{a.e.\ in~$Q$}},$ (3.29)
@@
as the functions $M_{\Omega}$, $M_{\Gamma}$, $\nu$, $\lambda$ and $\gamma$ are continuous. Since, due to (A2) and (A3), these functions are also bounded, we use Lebesgue’s dominated convergence theorem along with the weak convergences in (3.25) to infer that
$\nu(\varphi_{k})\mathbf{D}{\boldsymbol{v}}_{k}$ $\to\nu(\varphi)\mathbf{D}{\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d\times d})$},$ (3.31)
$\lambda(\varphi_{k}){\boldsymbol{v}}_{k}$ $\to\lambda(\varphi){\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.32)
$M_{\Omega}(\varphi_{k})\nabla\mu_{k}$ $\to M_{\Omega}(\varphi)\nabla\mu$ $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.33)
$\gamma(\psi_{k}){\boldsymbol{v}}_{k}$ $\to\gamma(\psi){\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$},$ (3.34)
$M_{\Gamma}(\psi_{k})\nabla_{\Gamma}\theta_{k}$ $\to M_{\Gamma}(\psi)\nabla_{\Gamma}\theta$ $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$}.$ (3.35)
Combining the convergences (3.25a)–(3.25f), (3.26)–(3.28) and (3.31)–(3.35), it is straightforward to pass to the limit as $k\to\infty$ in the discretized weak formulation (3.8) to conclude that the quintuplet $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ fulfills the variational formulation (2.16) for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$,
@@
Hence, because of density, (2.16) holds true for all test functions ${\boldsymbol{w}}\in{\bf V}_{\sigma,{\bf n}}$, $\zeta\in V$, ${\zeta_{\Gamma}}\in V_{\Gamma}$ and $(\eta,{\eta_{\Gamma}})\in{\cal V}={\cal V}_{K}$. This verifies Definition 2.4(ii).
@@
as the orthogonal projections converge strongly in $H$ and in ${H}_{\Gamma}$, respectively. On the other hand, it follows from the strong convergences in (3.25) and (3.25) that
@@
We still need to establish the weak energy dissipation law. Therefore, let $\rho\in C^{\infty}([0,T])$ be an arbitrary nonnegative test function. Employing the convergences (3.25) and (3.25), the weak lower semicontinuity of the mappings
@@
$\sqrt{\nu(\varphi_{k})}\,\mathbf{D}{\boldsymbol{v}}_{k}$ $\to\sqrt{\nu(\varphi)}\,\mathbf{D}{\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d\times d})$},$ (3.37)
$\sqrt{\lambda(\varphi_{k})}\,{\boldsymbol{v}}_{k}$ $\to\sqrt{\lambda(\varphi)}\,{\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.38)
$\sqrt{M_{\Omega}(\varphi_{k})}\,\nabla\mu_{k}$ $\to\sqrt{M_{\Omega}(\varphi)}\,\nabla\mu$ $\quad\text{weakly in $L^{2}(Q;{\mathbb{R}}^{d})$},$ (3.39)
$\sqrt{\gamma(\psi_{k})}\,{\boldsymbol{v}}_{k}$ $\to\sqrt{\gamma(\psi)}\,{\boldsymbol{v}}$ $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$},$ (3.40)
$\sqrt{M_{\Gamma}(\psi_{k})}\,\nabla_{\Gamma}\theta_{k}$ $\to\sqrt{M_{\Gamma}(\psi)}\,\nabla_{\Gamma}\theta$ $\quad\text{weakly in $L^{2}(\Sigma;{\mathbb{R}}^{d})$}.$ (3.41)
@@
Here, invoking the growth conditions from (R1), the equality in the last line follows by means of Lebesgue’s general convergence theorem (see [11, Section 3.25]) since the orthogonal projections in (3.9) converge strongly in $V$ and in ${V}_{\Gamma}$, respectively. As the nonnegative test function $\rho$ was arbitrary, this proves that the weak energy dissipation law stated in ((iv)) holds for almost all $t\in[0,T]$. As the time integral in this inequality is continuous with respect to $t$ and since the mapping $t\mapsto E_{K}\big{(}\varphi(t),\psi(t)\big{)}$ is lower semicontinuous, we conclude that ((iv)) actually holds true for all $t\in[0,T]$. This means that Definition 2.4(iv) is verified.
@@ 3.1.5 Additional regularity for the phase-fields
It remains to prove the additional regularities. Without loss of generality, we merely consider the case $d=3$ as the case $d=2$ can be handled analogously but is even easier due to the better Sobolev embeddings in two dimensions. We deduce from (2.16d) written for the solution $({\boldsymbol{v}},\varphi,\mu,\psi,\theta)$ that there exists a null set $\mathcal{N}\subset[0,T]$ such that
$\int_{\Omega}\nabla\varphi(t)\cdot\nabla\eta+\int_{\Gamma}\nabla_{\Gamma}\psi(t)\cdot\nabla_{\Gamma}{\eta_{\Gamma}}+\frac{1}{K}\int_{\Gamma}\big{(}\psi(t)-\varphi(t)\big{)}({\eta_{\Gamma}}-\eta)$
$\quad=\int_{\Omega}\Big{(}\mu(t)-F^{\prime}\big{(}\varphi(t)\big{)}\Big{)}\,\eta+\int_{\Gamma}\Big{(}\theta(t)-G^{\prime}\big{(}\psi(t)\big{)}\Big{)}\,{\eta_{\Gamma}}$ (3.42)
holds for all $t\in[0,T]\setminus\mathcal{N}$ and all test functions $(\eta,{\eta_{\Gamma}})\in{\cal V}_{0}$.
Let now $t\in[0,T]\setminus\mathcal{N}$ be arbitrary. We infer from (3.1.5) that the pair $\big{(}\varphi(t),\psi(t)\big{)}$ is a weak solution of the bulk-surface elliptic problem
$-\Delta\varphi(t)$ $=f(t)$ $\quad\text{in $\Omega$},$ (3.43a)
$-\Delta_{\Gamma}\psi(t)+\partial_{\bf n}\varphi(t)$ $=g(t)$ $\quad\text{on $\Gamma$},$ (3.43b)
$K\partial_{\bf n}\varphi(t)$ $=\psi(t)-\varphi(t)$ $\quad\text{on $\Gamma$},$ (3.43c)
@@
$f(t):=\mu(t)-F^{\prime}\big{(}\varphi(t)\big{)}\quad\text{and}\quad g(t):=\theta(t)-G^{\prime}\big{(}\psi(t)\big{)}.$
For the definition of a weak solution to such bulk-surface elliptic problems, we refer to [61, Definition 3.1].
$\left\|F^{\prime}\big{(}\varphi(t)\big{)}\right\|_{H}$ $\leq C+C\left\|\varphi\right\|_{L^{6}(\Omega)}^{3}\leq C,$ (3.44)
$\left\|G^{\prime}\big{(}\psi(t)\big{)}\right\|_{{H}_{\Gamma}}$ $\leq C+C\left\|\psi\right\|_{L^{2(q-1)}(\Gamma)}^{q-1}\leq C.$ (3.45)